Consider the matrix


A = 1 4

2 3


(a) Compute A-1

(b) Find det(A-1)

(c) Deduce a relation (if it exists) between det(A) and det(A-1)





1
Expert's answer
2021-05-13T13:32:42-0400

Solution.


"A=\begin{pmatrix}
 1 & 4\\
 2 & 3
\end{pmatrix}"

(a)

"det A=3-8=-5."

"A^{-1}=-\frac{1}{5}\begin{pmatrix}
 3 & -2 \\
 -4 & 1
\end{pmatrix}^T=-\frac{1}{5}\begin{pmatrix}
 3 & -4 \\
 -2& 1
\end{pmatrix}=\begin{pmatrix}
 -0.6 & 0.8 \\
 0.4 & -0.2
\end{pmatrix}."

(b)

"det (A^{-1})=0.12-0.32=-0.2=-\frac{1}{5}."

"det(A^{-1})=\frac{1}{det A}=-\frac{1}{5}."

(c)

"det(A^{-1})=\frac{1}{det A}."